OpenStudy (anonymous):

x^3-8=0 solve using an appropriate technique????

OpenStudy (kaiz122):

x^3=2^3

OpenStudy (kaiz122):

to get the value of x, simply find the cube root of booth sides of the equation

OpenStudy (anonymous):

so it would be 2 and 8 ?

OpenStudy (kaiz122):

no, \[x^3 =2^3\] \[\sqrt[3]{x^3}=\sqrt[3]{2^3}\] what would be x?

OpenStudy (anonymous):

2 and -2

OpenStudy (kaiz122):

in finding the cube root,there will only be one answer and it's positive.

OpenStudy (anonymous):

well i thought that too and on my test i put 2 and she said i needed another answer????

OpenStudy (kaiz122):

it's positive if the radicand is positive and negative if the radicand is negative

OpenStudy (kaiz122):

i think it's the only answer.

OpenStudy (anonymous):

ok thank you very much i will question her today on that issue.

OpenStudy (kaiz122):

sorry, but there are other ways for solving it. x³ - 8 = 0 (x - 2)(x² + 2x + 4) = 0 (x - 2) = 0 x = 2 x² + 2x + 4 = 0 x = {-2 ± √[2² - 4(1)(4)]}/2(1) x = {-2 ± √[4 - 16]}/2 x = {-2 ± √-12}/2 x = {-2 ± √[(4)(-3)]}/2 x = {-2 ± 2i√3]}/2 x = -1 ± i√3